0 0
Add a comment Improve this question Transcribed image text
Answer #1

Ennery of the system is

E = (1/2) KA^2 = (1/2)(223) (0.0460 M)^2 = 0.235 j

maximum speed of the object is

v_max= omega A = square root k/m( A)= square root 223/0.576 kg (0.0460 m) = 0.90 m/s

maximum accleration is

a_max= omega^2 A = (k/m) A = 223/0.576 kg (0.0460 m)= 17.8 m/s^2
Answer #2


here w = sqrt(k/m) = sqrt(223/0.576) = 19.67

A) E = 0.5*m*w^2*A^2 = 0.5*0.576*19.67^2*0.046^2 = 0.235 J..

B) Vmax = A*w = 0.046*19.67 = 0.90482 m/sec..

C) amax = A*w^2 = 0.046*19.67^2 = 17.79 m/s^2